OpenStudy (anonymous):

Rationalize the denominator of (sqrrt of 50)/(sqrrt of 5). Simplify.

OpenStudy (anonymous):

First, we should simplify it a bit: \[\sqrt{50}/\sqrt{5} = \sqrt{25*2}/\sqrt{5} = 5\sqrt{2}/\sqrt{5}\] now lets multiply the top and bottom by root 5 \[5\sqrt{2}/\sqrt{5}*(\sqrt{5}/\sqrt{5}) = 5\sqrt{10}/5\] the 5 on the top and 5 on the bottom cancel out and you get:\[\sqrt{10}\]

OpenStudy (anonymous):

Thank You :)